OpenStudy (anonymous):

check √(2x+5) for extraneous solutions and show your work.

OpenStudy (anonymous):

is that the full question?

OpenStudy (anonymous):

sorry the question is:check √(2x+5)=x+3 for extraneous solutions and show your work.

OpenStudy (anonymous):

okay i was about to say...

OpenStudy (anonymous):

lol I know

OpenStudy (anonymous):

well just solve for x, there should be 2 values then test your values for x and see if they both actually work

OpenStudy (anonymous):

when I solved for x I got (x+2)(x+2)

OpenStudy (anonymous):

which is basically just -2 rite?

OpenStudy (anonymous):

so....

OpenStudy (anonymous):

start by squaring everything to get rid of the radical 2x+5 = (x+3)^2 2x+5 = x^2 +6x +9 0 = x^2 +4x +4 (x+2)(x+2) x=-2 now lets test it √(2x+5)=x+3 √(-4+5)=-2+3 1 = 1 works no extraneous solutions

OpenStudy (anonymous):

ok that's what I had! but since I didn't find any I just wanted a 2nd opinion

OpenStudy (anonymous):

thank u! :)

OpenStudy (anonymous):

well it says just to check for them haha you're welcome